# Fully Distributed Fog Load Balancing with Multi-Agent Reinforcement Learning

### v1

@@ V Performance Evaluation
This paper mainly focuses on comparing the learning of independent decision-making policies with the learning of a common policy. A common policy exists in both single-agent and MARL CTDE solutions. Hence, we evaluate our proposed fully distributed MARL approach against a common agent policy as in [2]. This highlights the advantages of training independent agents compared to a single LB agent, including autonomy, adaptability, and scalability. For instance, $AP_{2}$ in Fig. 6 can quickly learn, after a few interactions, that $F_{1}$ is not feasible to process any of its generated workloads, given its distance and limited resources. Similarly, $AP_{1}$ can quickly learn that it can use $F_{2}$ to process its compute-intensive workloads, despite its distance.
Figure 6: Benefit of independent distributed policies.
Using a common policy to learn these infeasible decisions requires more interactions than independent policies, especially in large and complex environments. A common policy explores a larger number of state-action combinations, while independent policies provide more targeted exploration for each agent by focusing on the infeasible decisions for every agent separately. Independent policies mitigate the burden of exploring vast state and action spaces, allowing for faster learning even with larger spaces. Hence, the results of this section will help show the ability of our fully independent agents to learn better distribution strategies compared to a common centralized policy.
We leverage the system parameters based on our previous contributions [18, 17, 2], which ensures consistency and relevance in our evaluation, and provides valuable insights into the effectiveness and applicability of the fully distributed MARL approach in real-world scenarios. We used TF-Agents [25] to implement the DDQL algorithm with the hyper-parameters in Table III. A fully independent agent is deployed in each AP with its policy model, Q-Network, and experience replay buffer without sharing learning information between the agents. The simulations run on a DELL G5 laptop with 16GB RAM and Intel Core i7-9750H (6 physical cores, 12MB Cache, running up to 4.5 GHz). The agents are trained using NVIDIA GeForce RTX 2070 GPU with Max-Q Design (8GB RAM). To simulate realistic Fog environments we used a Discrete-Event Simulator called YAFS [26].
@@
Flat interconnected Fog topologies are generated using the Networkx library [27] as random undirected graphs resembling the Internet Autonomous System network. Nodes in the graph are then identified as Cloud, Fog nodes, or APs based on their betweenness centrality [28]. For instance, Fig. 7 shows a sample topology with a Cloud, 11 Fog nodes, and 21 APs. Fog nodes are then assigned heterogeneous compute resources, in Instructions Per Second (IPS), from evenly spaced values in the range $[10^{3},10^{5}]$. This is done inversely proportional to the number of directly connected APs to resemble unbalanced load distribution.
@@
Figure 7: Unbalanced flat Fog topology with interconnected Fog nodes.
@@
The Cloud, with $10^{6}$ computing resources, is connected to two Fog nodes with the highest betweenness. Fog nodes are assumed closer to IoT devices with low link latency compared to the Cloud, which incurs higher delays due to longer transmission distances and additional processing layers. To simulate the physical distances of IoT-Fog, Fog-Fog, and Fog-Cloud links, the propagation delay of those links, in seconds, is drawn from a uniform distribution in the ranges $[1,2)$, $[2,4)$, and $[10,20)$, respectively [7]. While their bandwidth, in Mbps, is drawn from a uniform distribution in the ranges $[10^{2},10^{3})$, $[10^{3},10^{4})$, and $[10^{3},10^{4})$, respectively [7]. Heavy, moderate, and light workloads require $10^{4}$, $10^{3}$, and $10^{2}$ compute instructions, respectively. The data sizes of these workload categories are $10^{3}$, $10^{2}$, and $10^{1}$ Bytes, respectively. These values were chosen relative to Fog and link resources to create a system with scarce resources relative to the demand.
@@
Workloads are generated as a Poisson Point Process using exponential distributions to effectively model the stochastic nature of job arrivals in Fog computing environments [29]. We used three scales for the exponential distributions, i.e., $\beta=$ 100, 150, and 200, to evaluate the performance under various load intensities, i.e., high, medium, and low generation rates/frequencies, respectively. Without efficient load distribution, the high rate can easily create a system bottleneck with heavy workloads and the unbalanced structure of the environment. In episodes of 10K simulation steps, training starts with the low rate for 30K training steps. A training step spans four decision steps, each spanning a variable number of simulation steps, depending on the workload generation rate. After 30K training steps, the rate instantly increases to medium, and then to high, requiring the agents to adapt to this rapid change using transfer learning.
@@
To evaluate the agents’ ability to reduce the number of queued jobs in each Fog node, we use the average system waiting delay as the main performance metric. We then evaluate balance and fairness by examining the similarity in waiting times and resource utilization across Fog nodes. Finally, we evaluate the system average for end-to-end workload execution delay, which includes network latency besides Fog waiting and service times. To underscore the benefits of learning independent policies over a common decision-making policy, we compare the proposed fully Distributed RL (DRL) approach against a Centralized single-agent RL (CRL) solution. We also include four baseline methods for completeness to show the significant difference between simple and intelligent load distribution solutions. These baselines include the selection of a random Fog node for each generated workload, and a Distributed Round-Robin (DRR) mechanism, where each AP assigns each workload among Fog nodes in a circular order. We also include the selection of the nearest Fog node to each workload’s AP. The final baseline is selecting the fastest Fog node for each workload based on compute requirements relative to node and link resources.
@@ V-A Performance Analysis
The results presented in this section are averaged over 10 experiment trials with different random seeds to showcase the confidence intervals across different Fog topologies generated with the random graph generator. Using confidence intervals provides a measure of the variability in the results to assess the precision and reliability of the average values obtained from multiple experiment trials, enhancing the robustness of the findings. Due to the significant performance differences between the proposed DRL approach and the simple baseline methods, we used a logarithmic scale to show the values of DRL in the figures that compare it with baselines. The results are gathered using the highest generation rate, After training on the three rates each with 30K training steps, as it is the most difficult for the agent to learn [2]. We use an evaluation episode of 100K simulation steps, i.e., 10×10$\times$10 × longer than training episodes to offer a more reliable measure of RL effectiveness and generalization capability in real-world scenarios.
Starting with the average system waiting delay, i.e., the average waiting time in the queues of all Fog nodes, Fig. 8LABEL:sub@fig:avgWaiting1 illustrates the performance of RL approaches compared to the baseline methods. In our evaluations, CRL performs better in reducing the system waiting delay than traditional baseline methods. This highlights the effectiveness of leveraging RL-based interactive learning to optimize load distribution in Fog environments. Furthermore, transitioning to DRL yields notable performance improvements beyond the CRL. By distributing decision-making across APs, the fully distributed independent agents capitalize on localized information and dynamic adjustments to the actions of other agents, resulting in even lower system waiting delays.
(a) DRL compared to CRL and baseline methods
@@
Fig. 8LABEL:sub@fig:avgWaiting2 focuses on the performance of both centralized and distributed RL solutions with respect to the frequency of observing the environment. It shows the performance of realistic observations using a Gossip protocol with 3 seconds interval between two consecutive actions. This is compared against the unrealistic assumption, commonly assumed in the literature, of readily available observations for every action. The figure highlights that realistic observations deliver slightly inferior performance compared to the optimal scenario where observations are unrealistically provided in real-time. This underscores the trade-off between realism and performance in RL-based load distribution solutions, where performance is directly related to the observation frequency. Despite the marginal performance degradation, realistic observation intervals align with practical constraints often encountered in real-world environments.
After evaluating the average waiting delay in the system, we show the variations in waiting delay among all Fog nodes, where lower values represent similar waiting times across the Fog nodes. Fig. 9LABEL:sub@fig:stdWaiting1 shows that RL-based solutions exhibit fewer variations than the baselines, with significantly better performance for DRL. By distributing RL decision-making, the agent in each AP manages the distribution of its workloads independently from other agents. This leads to a more effective load distribution considering more localized resources and adapting to the decisions of the other agents in the system. In contrast, CRL may exhibit higher variations in waiting time compared to DRL due to the centralized decision-making entity. The centralized agent struggles to efficiently adapt its decisions to the diverse workload demands from multiple APs across the heterogeneous Fog nodes.
(a) DRL compared to CRL and baseline methods
@@
Comparing DRL and CRL with respect to the frequency of observing the environment, Fig. 9LABEL:sub@fig:stdWaiting2 shows a slightly different trend when it comes to the variations in waiting delay between Fog nodes. DRL with real-time observations is the best to equalize the waiting time across all Fog nodes. Using interval-based observations, DRL still performs better than CRL but with more variability across the different trials. CRL with real-time observations, however, shows the largest variability across trials, indicating less confidence,i.e., uncertainty, on the behavior of the agents during deployment. Interestingly, CRL with interval-based observations, with slightly less equalized waiting delay across Fog nodes, has less variability across trials than real-time observations.
Fig. 10LABEL:sub@fig:stdUtilization1 shows DRL’s ability to provide fair resource utilization between Fog nodes by reducing the variation in Fog resource utilization across the network. It shows significant improvement in DRL resource utilization fairness compared to CRL and other baselines. This mitigates the probability of overloading or underloading individual Fog nodes in the system. Maximizing the utilization of available resources enhances system performance by encouraging all Fog nodes to contribute equally to the overall workload processing, relative to their capacity constraints. On the other hand, mitigating overloading enhances system stability by ensuring the maximum availability of spare resources across the network, which minimizes the probability of dropping workloads in overloaded Fog nodes in case of a sudden surge in workload generation.
(a) DRL compared to CRL and baseline methods
@@
Fig. 10LABEL:sub@fig:stdUtilization2 shows a small degradation in DRL fairness after using realistic interval-based observations. On the other hand, the choice between real-time and interval-based observations does not seem to have a significant difference in CRL fairness. Unlike Fig. 9LABEL:sub@fig:stdWaiting2, Fig. 10LABEL:sub@fig:stdUtilization2 shows small variability across the different trials, indicating less confidence on agents behavior. It is important to analyze agents’ ability to provide fairness in resource utilization, similar waiting delay between all Fog nodes, and minimal average waiting time. However, we still need to observe the average end-to-end execution delay, including network links latency, Fog queue waiting time, and Fog service time for all generated workloads.
Fig. 11LABEL:sub@fig:execution1 shows the ability of RL-based solutions to significantly minimize, compared to baseline methods, the average end-to-end execution delay for all generated workloads in the system. This performance improvement is directly attributed to agents’ ability to reduce the accumulation of workloads in the queues of Fog nodes, emphasizing the critical role of reducing the waiting delay in enhancing overall system performance. As shown earlier, there is a significant difference between DRL and CRL’s ability to minimize the variations in waiting delay and resource utilization between Fog nodes. However, the total execution delay achieved by both methods is quite similar, which is again attributed to their ability to reduce the waiting delay to a similar level relative to the performance of baseline methods.
(a) DRL compared to CRL and baseline methods
@@
The relatively small difference in total execution delay between DRL and CRL (see Fig. 11LABEL:sub@fig:execution2) underscores the efficacy of both approaches in optimizing system performance through different load distribution strategies. However, performance difference simply amplifies with less efficient solutions over longer simulations due to the accumulation of workloads over time, leading to increased total execution delays with DRL compared to CRL. Fig. 11LABEL:sub@fig:execution2 also shows a small degradation in DRL performance using interval-based observations compared to CRL, but still performance better than CRL. The performance of CRL, on the other hand, is almost identical between real-time and interval-based observations.
In summary, the results presented in this section offer insights into the performance of our proposed fully distributed MARL solution, i.e., DRL, against the performance of a single-agent RL solution and some other baselines. The results demonstrate the superior performance of RL approaches over baseline methods, with DRL showcasing the most notable improvements, highlighting the effectiveness of distributed decision-making in optimizing workload distribution. These improvements will amplify over longer simulations due to the accumulation of workloads over time, making such intelligent solutions the best option for real-world implementations. However, it’s crucial to note that the realistic performance of such solutions highly depends on observation frequency through multi-casting protocols, such as Gossip.

### v2

@@ V Performance Evaluation
This paper compares learning independent decision-making policies with learning a common policy, where a common policy results from single-agent RL, CTDE MARL, or even distributed MARL training with parameter or experience sharing. Hence, we compare the performance of learning independent policies by our proposed fully distributed MARL approach against learning a common policy by a single agent as in [2]. This comparison will highlight the advantages of training independent agents compared to a single LB agent, especially with convergence speed, performance, and scalability. To show why learning independent policies is faster than learning a common policy, Fig. 6 shows a scenario where considering the remote rich resources of Fog node $F_{2}$ is beneficial to the agent in $AP_{1}$ while considering the remote limited resources of Fog node $F_{1}$ is bad for the agent in $AP_{2}$. Both agents should learn to use nearby Fog resources while carefully considering only feasible remote Fog resources. Using a common policy to learn a different strategy for each $AP$ is harder, more complex, and requires more training than learning an independent policy for each $AP$. Learning a policy for each agent independently allows every agent to find its optimal load distribution strategy faster by avoiding the complexity of learning a global policy that satisfies the needs of all $APs$.
Figure 6: Learning a common policy vs. independently learned policies.
The system parameters of the experiments in this work are based on the experiments in our previous contributions [18, 17, 2], ensuring consistency and relevance in our evaluation. We used TF-Agents [25] to implement the DDQL algorithm with the hyper-parameters in Table III. A fully independent agent is deployed in each AP with its policy model, Q-Network, and experience replay buffer without sharing any information between the agents. The simulations run on a DELL G5 laptop with 16GB RAM and Intel Core i7-9750H (6 physical cores, 12MB Cache, running up to 4.5 GHz). The agents are trained using NVIDIA GeForce RTX 2070 GPU with Max-Q Design (8GB RAM). To simulate realistic Fog environments we used YAFS Discrete-Event Simulator [26].
@@
Our distributed agents are evaluated on isolated Fog regions randomly generated with complex flat topologies, i.e., multiple interconnected Fog nodes. To realistically simulate Fog networks, each Fog region is created, using the Networkx library [27], from random undirected graphs resembling Internet Autonomous System networks. The nodes in these graphs are identified as the Cloud, Fog nodes, or IoT APs based on their betweenness centrality [28], i.e., the Cloud is in the center of the graph while APs are at the edges (see Fig. 7). To resemble unbalanced load distribution relative to Fog resources, Fog nodes are assigned heterogeneous compute resources (uniformly between $[10^{3},10^{5}]$ Instructions Per Second (IPS)) inversely proportional to the number of their directly connected APs.
@@
Figure 7: A Fog region with unbalanced flat topology with 11 interconnected Fog nodes and 21 IoT APs.
@@
The Cloud is connected to two Fog nodes with the highest betweenness centrality and has computing resources of $10^{6}$ IPS. To simulate undedicated public networks with a distanced Cloud and Fog nodes towards the edge of the network, the propagation delays of IoT-Fog, Fog-Fog, and Fog-Cloud links are drawn from uniform distributions in the ranges $[1,2)$, $[2,4)$, and $[10,20)$ seconds, respectively [7]. The bandwidths of those link categories are drawn from uniform distributions in the ranges $[10^{2},10^{3})$, $[10^{3},10^{4})$, and $[10^{3},10^{4})$ Mbps, respectively [7]. To create a system with scarce resources, workloads with heavy, moderate, and light demands are simultaneously generated with compute requirements ($10^{4}$, $10^{3}$, and $10^{2}$ instructions, respectively) and data sizes ($10^{3}$, $10^{2}$, and $10^{1}$ Bytes, respectively) relative to Fog and link resources.
@@
These workloads are generated as a Poisson Point Process using exponential distributions to effectively model the stochastic nature of job arrivals in Fog computing environments [29]. We used three scales for the exponential distributions, i.e., $\beta=$ 200, 150, and 100, to evaluate the performance under low, medium, and high load intensities, respectively. Without efficient load distribution, heavy workloads with high generation rates can easily create system bottlenecks due to the unbalanced Fog environment. This makes learning optimal load distribution strategies with high generation rates much harder than learning those strategies in environments with less demand. Therefore, our agents are trained with a low rate first for 30K training steps (in episodes of 10K simulation steps) before the rate instantly increases to the next load intensity. Using transfer learning, the agents rapidly adapt to the increased generation rates as they intensify from low to medium, and then from medium to high.
@@
In the following section, we compare the proposed fully Distributed RL (DRL) approach against a Centralized single-agent RL (CRL) solution to underscore the benefits of independent decision-making over learning a common policy. To demonstrate the effectiveness of intelligent load distribution compared to simple selection methods, we also included Random and Distributed Round-Robin (DRR) selection, as well as selecting the Nearest Fog node to each AP and the Fastest Fog service for each workload. The average system waiting delay is the performance metric that allows us to evaluate the ability to reduce workload accumulation in each Fog node. We also plot the variations in waiting times and resource utilization between all Fog nodes in the region, which helps evaluate the agents’ ability to balance the load in the network and provide fairness in resource utilization, respectively. We then measure the average end-to-end execution delay for the IoT application loop to analyze the impact of network latency and service times on load distribution decisions based solely on reducing waiting delay.
@@ V-A Performance Analysis
The results in this section are averaged over 10 experimental trials, each with a different random seed, to illustrate the confidence intervals of performance across random variations in Fog architecture. These confidence intervals capture the variability in results due to topology changes, aiding in assessing our findings’ precision, reliability, and robustness across different environments. These results represent the performance in an evaluation episode that is 10×10$\times$10 × longer than training episodes, i.e., 100K simulation steps, offering a measure of generalization for lifelong inference during deployment. The workload generation rate in this evaluation episode is high to evaluate load distribution with frequently generated workloads, which is very difficult to learn [2]. When compared against baseline methods, a logarithmic scale is used on the results to show the performance of RL-based solutions due to their significantly superior performance.
Fig. 8LABEL:sub@fig:avgWaiting1 illustrates the mean system waiting delay achieved using each load distribution method, which is the average time a workload must wait in the queue of a Fog node before being served. The results show superior performance with RL-based methods compared to simple baselines, highlighting the need for intelligent load distribution in unbalanced Fog environments. The results also show notable performance gains when transitioning from centralized to distributed RL, achieved by learning independent load distribution strategies that leverage resources in proximity differently for each distributed agent. Blindly learning separate policies by independent agents not only outperforms learning a common policy for all APs, but also eliminates the need to share network parameters, buffered interactions, or the actions of other agents. The small error bars in the figure show the robustness of RL-based solutions against variations in Fog topology across multiple experimental runs, showing the learning ability of the agents despite changes in problem complexity.
(a) DRL vs. CRL and baselines
@@
Fig. 8LABEL:sub@fig:avgWaiting2 compares DRL with CRL only focusing on the impact of practical interval-based observations (observing Fog load every 3 seconds) compared to unrealistic real-time observations (a new observation for every action). This comparison is essential to show the trade-off between theoretical performance and realism of practical deployment constraints. This trade-off is often neglected in the literature by assuming Fog load information (from across the network/region) is readily available before every action. The figure shows that using the same observation to make multiple assignment decisions has a larger impact on DRL than CRL. The slightly inferior performance of DRL with interval-based observations, compared to the unrealistic optimal solution where observations are unrealistically provided in real-time, is due to the non-stationarity problem created by outdated information observed by independent agents. The impact on CRL with interval-based observations is smaller than DRL because the non-stationarity does not exist with CRL as the new observation depends on the behavior of a single agent.
Fig. 9LABEL:sub@fig:stdWaiting1 shows the variations in waiting delays between Fog nodes, measured using the standard deviation of waiting times across all Fog nodes. Lower values represent more equalized or balanced waiting times between Fog nodes. The error bars in the figure indicate the variability in these values across different experiments, demonstrating the level of result consistency under various Fog topologies. The results show how intelligent RL-based solutions can better balance the waiting times across Fog nodes than simple baselines, proving their ability to learn efficient load distribution strategies that minimize workload accumulation in all Fog nodes. By distributing RL decision-making, each agent can independently learn to distribute its load by leveraging nearby resources based on its location in the network. Therefore, DRL can balance the waiting times better than CRL since independently learning effective load distribution strategies, that fit the needs of each agent, is faster than learning a common policy for all APs.
(a) DRL vs. CRL and baselines
@@
To focus on the impact of interval-based observations on balancing the waiting times between Fog nodes, we again limit our comparison in Fig. 9LABEL:sub@fig:stdWaiting2 to DRL and CRL only. The figure confirms that due to interval-based observations, outdated observations have a negligible impact on CRL compared to DRL. DRL is more affected by outdated observations because of the environment’s non-stationarity caused by the unobserved actions of multiple independent agents. With CRL, a single agent alters the state of the environment, making it stationary from that agent’s perspective, even if the same observation is used for multiple actions. However, DRL’s ability to balance waiting times between Fog nodes remains significantly better than CRL’s, with real-time and interval-based observations. With real-time observations, DRL demonstrates an impressive ability to achieve theoretically optimal load distribution, resulting in almost identical waiting times across all Fog nodes in the environment.
Besides balancing waiting times between Fog nodes, RL-based solutions balance the consumption of Fog resources by reducing the variation in resource utilization between Fog nodes (see Fig. 10LABEL:sub@fig:stdUtilization1). Balancing resource utilization provides fairness across Fog nodes and mitigates the probability of overloading or underloading individual Fog nodes in the region. This enhances system performance by encouraging all Fog nodes to equally contribute with their resources relative to their capacity and constraints. It also improves system stability by ensuring the maximum availability of spare Fog resources in the network, minimizing workload-dropping probability in case of a sudden surge in generation rates. While both RL methods are better than the baselines, DRL is significantly better in providing fair resource utilization using independently learned load distribution strategies. This proves the ability of our fully distributed agents to balance the system load without requiring the exchange of coordination information.
(a) DRL vs. CRL and baselines
@@
Once again, we narrow down the comparison to DRL and CRL only to study the impact of interval-based observations on balancing resource utilization between Fog nodes (see Fig. 10LABEL:sub@fig:stdUtilization2). The figure confirms our previous findings that the choice of observation frequency, i.e., real-time and interval-based observations, does not impact the performance of CRL due to having a single decision-making entity. Whether the CRL agent observes the environment before every action or makes an observation for all actions within 3 seconds, the variation in resource utilization (fairness) between Fog nodes is almost identical. DRL, however, is slightly impacted when observing the environment less frequently compared to unrealistically observing real-time Fog load before every action. Observing the load from Fog nodes once every 3 seconds is more practical for the actual implementation and deployment of the solution, and a slight degradation in the achievable performance can be tolerated given the constraints of real-world networks.
To analyze the impact of network latency and service times on load distribution decisions based solely on reducing waiting delay, we need to show the average end-to-end execution delay for the whole IoT application loop. This will demonstrate the direct relationship between minimizing the waiting delay of individual workloads and minimizing the execution delay of the flow of workloads that make IoT applications. Fig. 11LABEL:sub@fig:execution1 shows the ability of RL-based solutions to significantly minimize, compared to the simple baselines, the average execution time of the three application loops simultaneously running in the system. DRL performance is still slightly better than CRL due to the independent agents that learn different load distribution strategies based on their relative locations to Fog nodes, reducing workload accumulation in all Fog nodes better than using a single policy for the whole region. These independent agents find the optimal solution faster than learning a common policy by the single-agent solution. By avoiding the exchange of coordination information, and its associated overhead, these independent agents will also converge faster than implicitly learning a common policy using distributed agents that share their network parameters, experiences, or actions.
(a) DRL vs. CRL and baselines
@@
With Fig. 11LABEL:sub@fig:execution2, we can analyze the impact of interval-based observations on the average end-to-end execution delay of IoT application loops. Once again, the results show that real-time and interval-based observations lead to almost identical average execution delays using CRL. Even with outdated observations, the series of actions, by a single agent, between two consecutive observations determine the next observed state of the environment, leading to environment stationarity. Although DRL achieves a smaller average execution delay than CRL, DRL is slightly affected by these outdated observations due to having independent actors that simultaneously change the state of the environment. Even with interval-based observations, the performance of DRL is still better than CRL. The design of smaller regions can reduce DRL’s non-stationarity problem by minimizing the maximum communication delay for multi-casting Fog load information. This reduction enables the selection of smaller, yet practical, observation intervals based on the dimensionality of the region, lowering the impact of interval-based observations by independent decision makers.
In summary, the results presented in this section demonstrate the superior performance of our proposed fully distributed MARL solution (DRL) compared to the CRL approach (decision-making using a common policy) and other baselines. The results highlight the ability of independently trained agents to learn distinct load distribution strategies more quickly than when learning a common policy. These strategies collectively reduce workload accumulation in Fog nodes more effectively than a CRL policy with the same amount of training. The fully distributed agents effectively minimize average waiting delay, balance waiting times, and ensure fair resource utilization across all Fog nodes in their region, thereby improving the average end-to-end execution delay.
Without centralized or inter-agent coordination, our fully distributed agents can jointly learn optimal load distribution strategies more quickly and with less overhead than learning a common policy. A common policy is typically learned explicitly by a single agent or CTDE solutions, or implicitly through parameter, experience, or action sharing. Also, learning independent policies that optimize a common objective enables the management of smaller collaboration regions within a global network, while still ensuring global optimality. Using small collaboration regions allows for a scalable solution that reduces problem complexity since the agents have fewer Fog nodes to manage, i.e., state and action spaces are smaller, making learning easier and converging faster. The results also emphasize the trade-off between performance and real-world applicability, as we assess the impact of interval-based observations rather than assuming idealized conditions where Fog load information is continuously available before every action.